Solve xd^2y/dx^2-(2x-1)dy/dx+(x-1)y=0


1
Expert's answer
2021-06-18T12:40:49-0400

"x\frac{d^2y}{dx^2}-(2x-1)\frac{dy}{dx}
+(x-1)y=0\\
\frac{d^2y}{dx^2}-(2-\frac{1}{x})\frac{dy}{dx}+(1-\frac{1}{x})y=0\\
Here, p=-2+\frac{1}{x}, q=1-\frac{1}{x}\\
Then,\\
1+p+q=0 \implies y_1=e^x \, \text{is a solution of given differential equation.}\\
\text{
Now, put} \,y=vy_1=ve^x, \\ \text{
and corresponding }\,y' and \,y'' \text{
values in the given differential equation, we get}\\
\frac{d^2v}{dx^2}+{1}{x}\frac{dv}{dx}=0\\
Put\\
 u=\frac{dv}{dx}.\\\text{
Therefore, the above differential equation is} \,
\frac{du}{dx}+\frac{u}{x}=0\\\text{
Using method of separation of variables},\\
\frac{du}{u}=-\frac{dx}{x}\\
logu=-logx+logc_1\\
u=\frac{c_1}{x}\\
\frac{dv}{dx}=\frac{c_1}{x}\\
v=c_1logx+c_2\\
ye^{-x}=c_1logx+c_2\\
y(x)=e^{x}(c_1logx+c_2)\\
\text{This is the required solution.}"